### v3

# Training-Free Bias Mitigation by LLM-Assisted Bias Detection and Latent Variable Guidance

Training-Free Bias Mitigation by LLM-Assisted Bias Detection and Latent Variable Guidance
Jinya Sakurai,  Yuki Koyama,  Issei Sato
The University of Tokyo
sakurai-jinya725@g.ecc.u-tokyo.ac.jp,   koyama@pe.t.u-tokyo.ac.jp,
sato@g.ecc.u-tokyo.ac.jp
Corresponding author: sakurai-jinya725@g.ecc.u-tokyo.ac.jp
1 Introduction
Figure 1: Overview of FairT2I. The LLM detects biases that exist in the input prompt and converts it into a set of bias-aware prompts by sampling the detected attributes from fair distributions. The interactive interface allows users to inspect, edit, and regenerate images, enabling both societal bias mitigation and improved diversity.
Text-to-image (T2I) models, such as Stable Diffusion (sd3; sdxl), Imagen (imagen3), and Flux (flux), have transformed visual content creation by allowing users to generate high-quality images from natural language descriptions. Yet, despite their accessibility and creative potential, these models inherit and amplify the biases present in large-scale web datasets (bavalatti2025systematic; naik2023social). When prompted with neutral descriptions, models often produce stereotypical representations such as depicting a nurse as female or a chief executive officer as male, thereby reinforcing societal inequalities (wan2024survey; mandal2024generated; sandoval2024perpetuation; zhang2024partiality; dealmeida2024bias). The increasing reuse of synthetic images in training pipelines further intensifies these biases over time (wyllie2024fairness; alemohammad2023self).
Many algorithmic approaches have been proposed to mitigate this problem; however, they continue to face significant practical limitations from the user’s perspective. Large-scale dataset curation is challenging to sustain over time (bianchi2023easily), and most existing strategies (Kim_2024; zhang2023iti; hirota2024saner) rely on computationally expensive fine-tuning or retraining for each target bias, often involving delicate hyperparameter tuning (friedrich2023fair). In addition, these approaches typically depend on manually pre-defined attribute lists (Kim_2024; zhang2023iti; friedrich2023fair), which not only constrains their adaptability to emerging or context-dependent biases but also limits the degree of user control and interpretability during deployment.
To overcome these constraints, we present FairT2I, a novel, theoretically grounded framework that enables bias-mitigated diffusion-based T2I process. Figure 1 illustrates the overview. FairT2I combines the reasoning ability of large language models (LLMs) with a flexible bias-aware image generation process. FairT2I operates entirely at inference time without any model fine-tuning or pre-defined attributes, allowing users to directly intervene in the debiasing process. The framework consists of three main components:
  1. 1.
    LLM-based bias detection: Given a user prompt, the LLM analyzes the text to identify implicit categories and attributes that are likely to introduce bias in generated images, such as gender, ethnicity, or age. This transforms unstructured text into explicit bias dimensions that can be interpreted and controlled.
  2. 2.
    Attribute resampling: Once potential biases are identified, users can control the relative importance of each attribute (i.e., attribute distribution). The system supports multiple types of distributions, including a uniform distribution, one based on real-world statistics, and a custom user-specified distribution, enabling flexible control over fairness and diversity. Based on the selected distribution, the system resamples attributes to reflect the desired balance. This resampled distribution is then used to generate bias-aware prompts that explicitly encode the desired attribute composition.
  3. 3.
    Latent variable guidance: The image generation process is mathematically decomposed into attribute-conditioned components, and FairT2I reweights these components according to the adjusted attribute distribution. By leveraging the bias-aware prompts derived from attribute resampling, the framework realizes latent variable guidance in a way that is not ad hoc but theoretically grounded. This approach enables bias mitigation while maintaining fidelity and diversity in T2I, all without retraining the underlying diffusion model.
Interactive Interface for Human–AI Collaboration
To support transparency and user control, we designed a web-based interface that operationalizes the bias mitigation pipeline of FairT2I. Once users give an input prompt, the system automatically detects bias categories and displays them with associated attributes. Users can then add, remove, or modify attributes and adjust their weights through sliders if desired. After confirmation, FairT2I performs attribute resampling using the bias-adjusted attribute distribution and generates a set of images. This process allows users to recognize how textual biases influence visual outcomes and provides direct agency to guide the model toward fairness or greater diversity.
Contributions
The main contributions of this work are as follows.
  • We introduce FairT2I, a training-free framework that integrates LLM-based bias detection with a mathematically grounded latent variable guidance formulation that factorizes the generative process into attribute-conditioned components. This is more than a simple combination of an LLM and a T2I model—our principled formulation provides a mathematical interpretation of potential sources of societal bias within the T2I generation process, offering a unified perspective that can encompass a wide range of existing ad hoc debiasing techniques.
  • We design an interactive interface that allows users to visualize, modify, and control bias attributes in real time, bridging fairness research and human–AI collaboration.
  • Through automatic and human-in-the-loop evaluations, we demonstrate that FairT2I outperforms baseline models in both societal bias mitigation and image diversity, while maintaining image quality and prompt fidelity.
2 Related Work
Text-to-image (T2I) models
The generative capabilities of T2I models (imagen3; sd3; chen2025pixart; flux) have improved dramatically through several key developments: training on large scale text image pairs (schuhmann2022laion; Chen_2023), architectural improvements (peebles2023scalable) from UNet (ronneberger2015u) to transformer-based designs (AttentionAllYouNeed; Dosovitskiy_2020), and theoretical developments from diffusion models (sohl2015deep; ddpm; Song_2020) to flow matching (Lipman_2022; Xingchao_2022; lipman2024flowmatchingguidecode). Furthermore, recent fine tuning approaches have significantly reduced generation time by minimizing the required number of inference steps (liu2023instaflow; chen2024pixartdeltafastcontrollableimage; sauer2024fast; Sauer_2023; Yang_2023). A distinctive characteristic of state- of-the-art models is their use of separately trained components, such as Variational Auto Encoders (kingma_2014) for image compression, text encoders (raffel2020exploring; clip), and flow model backbones, each trained on different datasets.
Societal bias in T2I models
Numerous studies have investigated biases in T2I models. Research such as Ghosh_2023; Wu_2023; bianchi2023easily highlights the biases embedded in generated outputs for seemingly neutral input prompts that lack explicit identity- or demographic-related terms. Other works, including Cho_2022; Luccioni_2023 predefine sensitive human attributes and analyze biases in outputs generated from occupational input prompts. naik2023social provides broader analyses, including comparative studies with statistical data or image search results, as well as spatial analyses of generated images. Wang_2023 applies methods from social psychology to explore implicit and complex biases related to race and gender. Furthermore, Luccioni_2023 introduces an interactive bias analysis tool leveraging clustering methods. Lastly, Chen_2024 examines the potential for AI-generated images to perpetuate harmful feedback loops, amplifying biases in AI systems when used as training data for future models. D'Incà_2024; Chinchure_2023 employ LLMs to detect open-ended biases in text-to-image models where users do not have to provide predefined bias attributes.
Bias mitigation in T2I models
One line of work focuses on making T2I models fair by fine-tuning or retraining various components, from the diffusion backbone (Kim_2024) and text encoder hirota2024saner to postfix prompt embeddings zhang2023iti or additive residual image vectors Seth_2023_CVPR. While effective, such retraining is both time- and resource-intensive and must be repeated whenever societal notions of bias evolve. By contrast, several methods eliminate societal bias only by steering generation at inference time. FairDiffusion friedrich2023fair guides sampling with randomly chosen bias attributes; (chuang2023debiasing) removes unwanted directions from the text embedding space via calibrated projection matrices; and ENTIGEN bansal2022well simply appends an ethical fairness postfix to the input prompt. Although these approaches require no additional training, they do depend on a predefined list of bias categories and attribute values for control.
Interactive interface for T2I generation
Recent research has increasingly explored user interfaces that support interactive control and understanding of T2I generation. PromptCharm (wang2024promptcharm) introduced a mixed initiative interface that refines prompts through multi-modal feedback and visual explanations, allowing users to iteratively steer image generation with system suggestions. Promptify (brade2023promptify) similarly employed LLMs to recommend prompt modifications based on user feedback, enabling interactive exploration and refinement of text prompts. Beyond linguistic prompting, PromptPaint (chung2023promptpaint) introduced an interactive canvas that enables users to spatially compose scenes by assigning distinct text prompts to different regions of the image. PromptMap (adamkiewicz2025promptmap) shifted the focus from control to sensemaking, visualizing the branching structure of prompt evolution as a map to help users navigate and reflect on their creative process. Expandora (choi2025expandora) extended these ideas to structured exploration, allowing users to specify design intentions and desired diversity levels to automatically generate a wide range of prompt variations. Collectively, these systems emphasize human agency, transparency, and creativity in T2I interaction. In contrast, FairT2I focuses on societal bias detection and mitigation, offering an interface that reveals and allows direct manipulation of latent bias attributes, thereby bridging algorithmic fairness and interactive generative control.
3 Methodologies
In this section, we present FairT2I, a mathematically grounded framework for mitigating societal biases in text-to-image (T2I) generation. At its core, our approach introduces a latent variable guidance formulation that factorizes the generative process into attribute-conditioned components, making explicit how sensitive attributes influence image generation and enabling principled bias control within the model. Building upon this foundation, we incorporate (1) an LLM-based bias detection mechanism that infers bias-prone categories and attributes directly from input prompts without relying on predefined attribute sets, and (2) an attribute resampling module that redefines attribute distributions, for example according to real-world demographic statistics, to rebalance outputs and steer generation toward greater fairness.
3.1 Preliminaries: Classifier-Free Guidance
In diffusion models (ddpm; sohl2015deep) and flow matching models (sd3; Lipman_2022), classifier-free guidance (ho2022classifier) is a powerful tool for sampling images $\mathbf{x}$ conditioned on a text input $\mathbf{y}$. It accomplishes this by guiding an unconditional image generation model using the score function $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. This enables the sampling from
$p_{w}(\mathbf{x},\mathbf{y})\propto p(\mathbf{x})p(\mathbf{y}\mid\mathbf{x})^{w}\propto p(\mathbf{x})^{1-w}p(\mathbf{x}\mid\mathbf{y})^{w}$
where $w\in\mathbb{R}$ is the guidance scale typically used with $w>1$. The score satisfies
$\nabla_{\mathbf{x}}\log p_{w}(\mathbf{x},\mathbf{y})=(1-w)\nabla_{\mathbf{x}}\log p(\mathbf{x})+w\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{y})$
so by training the network to predict both the unconditional score $\nabla_{\mathbf{x}}\log p(\mathbf{x})$ and conditional score $\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{y})$, flexible sampling from the conditional distribution can be achieved through a weighted sum of the network outputs.
Whereas in classifier-free guidance, the weights for the score functions are typically set to achieve a trade-off between sample quality and text fidelity, our formulation 3 sets these weights to ensure fair image generation.
3.2 Latent Variable Guidance for Bias Mitigation
In latent variable guidance, societal bias is controlled by decomposing the network’s predicted score function into a weighted sum of component score functions, each conditioned on a distinct latent variable. These weights are then adjusted to enforce fairness in the generated outputs. Central to our formulation is the following proposition.
Proposition 1.
Let $\mathbf{y}$ represent the input text, $\mathbf{x}$ the generated image, and $\mathbf{z}$ a discrete latent variable taking values in a finite set $\mathcal{Z}$, and assume conditional independence between $\mathbf{x}$ and $\mathbf{z}$ given $\mathbf{y}$, the following equation holds:
$\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{y})=\sum_{z\in\mathcal{Z}}p(\mathbf{z}=z\mid\mathbf{y})\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y}).$ (1)
Proof.
See Appendix A. ∎
This formulation, presented in Equation (1), effectively decomposes the score function guiding the image generation process. It explicitly separates the influence of the latent attribute $\mathbf{z}$ (which is sampled according to its conditional probability $p(\mathbf{z}\mid\mathbf{y})$) from the subsequent image generation step, which is guided by both this sampled attribute $\mathbf{z}$ and the original input text $\mathbf{y}$. This perspective allows us to contextualize existing bias mitigation strategies. Approaches such as replacing text encoders with fair alternatives (hirota2024saner; chuang2023debiasing), learning fair prompt embeddings (zhang2023iti), or randomly sampling sensitive attributes (friedrich2023fair) can all be interpreted as methods that implicitly or explicitly substitute the T2I model’s original, potentially biased distribution $p_{\text{bias}}(\mathbf{z}\mid\mathbf{y})$ with a more equitable target distribution $p_{\text{fair}}(\mathbf{z}\mid\mathbf{y})$. Figure 2 illustrates the generation process with latent variable guidance, compared to the standard process without it, for an input prompt $\mathbf{y}=\textit{A CEO}$ and an attribute set $\mathcal{Z}=\{z_{1}=\textit{male},z_{2}=\textit{female}\}$. In the standard setting, images are denoised at each timestep under the biased distribution $p_{\mathrm{bias}}(\mathbf{z}\mid\mathbf{y})$, which tends to steer the generation toward male CEO representations. In contrast, latent variable guidance enables the replacement of $p_{\mathrm{bias}}(\mathbf{z}\mid\mathbf{y})$ with a fair distribution $p_{\mathrm{fair}}(\mathbf{z}\mid\mathbf{y})$, allowing the generation process to be steered toward more balanced and fair outcomes.
Figure 2: Comparison of generation processes without (left) and with (right) latent variable guidance for the input prompt $\mathbf{y}=\textit{A CEO}$ and attribute set $\mathcal{Z}=\{z_{1}=\textit{male},z_{2}=\textit{female}\}$. In the standard diffusion process, the model internalizes societal biases from the training data and denoises images toward male CEO representations, resulting in biased outputs. Latent variable guidance enables the adjustment of the denoising direction based on a fair attribute distribution, thereby allowing the mitigation of such biases in the generated results.
In practice, computing the summation over all possible values of $z$ in Equation (1) can be computationally prohibitive, especially when dealing with a large or continuous space of latent attributes. To overcome this computational challenge, we employ Monte Carlo sampling from the distribution $p(\mathbf{z}\mid\mathbf{y})$. Specifically, we approximate the expectation in Equation (1) using a finite number of samples. For the simplest case, using an approach with a single sample, this approximation becomes:
$\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{y})\approx\nabla_{\mathbf{x}}\log p\bigl(\mathbf{x}\mid\widetilde{\mathbf{z}},\mathbf{y}\bigr),\ \text{where}\ \widetilde{\mathbf{z}}\sim p(\mathbf{z}\mid\mathbf{y}).$
While this approach using a single sample provides an unbiased estimate of the true score, it may exhibit higher variance compared to the exact summation. This variance can be mitigated by using additional samples; however, this benefit comes at the cost of increased computation time. Consequently, a careful trade-off between computational efficiency and the accuracy of the score estimation must be considered when implementing this approach in practical scenarios.
3.3 LLM-Assisted Bias Detection
To implement latent variable guidance, we need to define a candidate set of biases $\mathcal{Z}$. The simplest approach is to predefine a closed set of biases, such as race and gender; however, this approach has several limitations. For instance, one significant limitation relates to computational challenges. The diversity of input prompts is virtually infinite, meaning that a predefined set of latent attributes $\mathcal{Z}$ can only appropriately handle a limited subset of these cases. Consequently, it is practically impossible to predefine a suitable $\mathcal{Z}$ for every conceivable prompt in advance. Another limitation concerns incomplete representation. If the latent attribute set $\mathcal{Z}$ is defined manually in a rule-based manner, it may fail to fully capture the diversity and context of the real world. This approach also risks overlooking biases embedded in the input text that are beyond human recognition.
To address these challenges, we leverage large language models (LLMs) (gpt4; anthropic2025claude37sonnet) to automatically detect biases in the input text, as with existing bias detection methods (D'Incà_2024; Chinchure_2023). Specifically, we use the LLM to predict the set of possible latent attributes $\mathcal{Z}$ from the input text $\mathbf{y}$. LLMs are prompted to output a set of latent attributes that are likely to appear in images generated by T2I models with the input text in a JSON format. This approach allows us to handle a broader range of input prompts and to detect biases that may not be apparent to human annotators.
3.4 Attribute Resampling
When we have a set of latent attributes $\mathcal{Z}$ from LLM as potential biases in the input prompts, we can formulate $p_{\mathrm{fair}}(\mathbf{z}=z\mid\mathbf{y})$ to perform sampling that mitigates bias. We consider two approaches to defining $p_{\mathrm{fair}}(\mathbf{z}=z\mid\mathbf{y})$.
Uniform distribution
One simple approach is to set the distribution of latent attributes to be uniform across all possible values of $z$: $p_{\mathrm{fair}}(\mathbf{z}=z\mid\mathbf{y})=\frac{1}{|\mathcal{Z}|}$, which corresponds to simply mixing the scores $\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y})$ with equal proportions across all possible values of $z$. This formulation coincides with that of Fair Diffusion (friedrich2023fair).
Employment statistics
Research has shown that T2I models tend to exaggerate demographic stereotypes beyond what we observe in real-world distributions across various latent attributes (naik2023social). One way to address this issue is to incorporate real-world statistical data as $p_{\mathrm{fair}}(\mathbf{z}=z\mid\mathbf{y})$, ensuring that the generated image distributions are at least as balanced as real-world demographics. [^1] It is worth noting that real-world occupational distributions often reflect systemic biases and unequal access to opportunities, shaped by historical and societal factors such as limited access to education or workplace discrimination. These disparities highlight that real-world distributions are not inherently fair.
Upon sampling the latent attribute $z$ from the fair conditional distribution $p_{\mathrm{fair}}(\mathbf{z}=z\mid\mathbf{y})$, we need to compute the score $\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y})$. To enforce that the T2I model incorporates the sampled attribute $z$, we use an LLM to naturally fuse the textual input $\mathbf{y}$ and the latent attribute $\mathbf{z}$, and feed this augmented text $\mathbf{y}\oplus\mathbf{z}$ directly into the T2I model in place of $\mathbf{y}$.
4 Evaluation of LLM-Detected Bias Categories
This section evaluates the bias categories and attributes surfaced by the LLM, using human-annotated lists as a reference.
4.1 Experimental Settings
Evaluation dataset
We used two datasets for this study: one for assessing societal biases in professions and another for examining potential biases in more diverse textual prompts. To evaluate occupational bias, we used the Stable Bias profession dataset (Luccioni_2023), which contains $131$ occupations collected from the U.S. Bureau of Labor Statistics (BLS) (see Appendix A of Luccioni_2023 for full details). To probe broader forms of bias, we adopt the Parti Prompt dataset (yu2022scaling), which includes more than $1,600$ English prompts designed to evaluate text-to-image generation models and test their limitations. For efficiency, we randomly sample $50$ prompts from each dataset, with the selected subset listed in subsection B.1.
LLMs for Bias Detection
We used Claude 3.7-sonnet for LLM-assisted bias detection, following a small pilot study that compared GPT-4o gpt4, Claude 3.7-sonnet anthropic2025claude37sonnet, Llama 3.3 grattafiori2024llama3herdmodels, and DeepSeek V3 liu2024deepseek. Further details of this pilot study are provided in the appendix Appendix B.
4.2 Human Annotation and Reference Set Construction
To establish a human reference for bias categories, we conducted an annotation study using the same prompts as in the evaluation datasets. 100 text prompts, consisting of 50 prompts from the Stable Bias dataset (Luccioni_2023) and $50$ prompts from the PartiPrompts dataset (yu2022scaling) were divided into five groups of $20$ prompts, with each group assigned to five independent annotators ($25$ participants in total). Annotators were instructed to list potential bias categories (e.g., gender, race, attire) along with multiple attributes per category (e.g., male, female, non-binary for gender). Clear format guidelines and illustrative examples were provided to ensure consistency. To maintain parity, the same few-shot examples were used both in the human instructions and in the system prompts provided to the LLMs.
For each prompt, we consolidated the five human responses by normalizing synonyms (e.g., ethnicity $\rightarrow$ race) and merging duplicates. The resulting union of all annotators produced a single consolidated human reference set of categories and attributes per prompt. Full details of the instructions, interface, and annotation guidelines are provided in subsection B.3.
Figure 3: Number of detected categories and attributes per category for the Stable Bias (left) and Parti Prompt (right) datasets. Bars show Human Mean (blue), Human Union (orange), and LLM (green), with error bars indicating one standard deviation.
4.3 Results and Discussion
Categories and attributes
To assess the ability to cover a broad range of bias categories and attributes, we compared the number of detected categories and attributes. Figure 3 presents results for the Stable Bias dataset (left) and the Parti Prompt dataset (right), showing the average of five annotators (human mean; blue), the consolidated union of all annotators (human union; orange), and the LLM outputs (green). On the Stable Bias dataset, the human union and the LLM produced a comparable breadth of categories ($9.18$ vs. $9.06$ on average), although the LLM generated more attributes per category ($4.54$ vs. $3.84$). On the Parti Prompt dataset, however, humans (union) enumerated more categories than the LLM ($10.52$ vs. $8.44$), while the LLM provided denser attribute lists ($5.59$ vs. $4.01$). Notably, the human mean values are consistently lower than both the union and the LLM (for example, only $2.76$ categories and $3.46$ attributes in Parti Prompt), reflecting high inter-annotator variability. Taken together, these findings suggest that, while humans provide greater breadth when aggregated, LLMs offer more consistent depth within categories.
Table 1: Gender and race attribute sets on the Stable Bias dataset. Few-shot examples show the categories provided in the annotation instructions and LLM prompt. Rows report the most frequent, minimum, and maximum sets for the human union and the LLM. Numbers in parentheses indicate the number of prompts in which the set occurred. For race, humans did not annotate the category in 4 prompts.
Gender Race
Few-shot examples non-binary, female, male White, Asian, Black, Hispanic
Human Union
   Most frequent non-binary, female, male (29/50) White, Asian, Black, Hispanic (16/50)
   Minimum non-binary, female, male None
   Maximum non-binary, other, male, les, gay, female Mixed, Asian, Hispanic, African-American, Arab, African, White, Native-American, Black
LLM
   Most frequent non-binary, female, male (46/50) Middle Eastern, Asian, Hispanic, White, Indigenous, Black (13/50)
   Minimum non-binary, female, male Middle Eastern, Asian, Hispanic, White, Black
   Maximum non-binary, female, androgynous, male Middle Eastern, Asian, White, Hispanic/Latino, South Asian, Indigenous, Black
Societal biases.
Table 1 illustrates the gender and race attribute sets detected on the Stable Bias dataset. On average, human consolidation yielded $3.52$ gender attributes and $5.34$ race attributes per prompt, while the LLM produced $3.08$ and $5.8$, respectively. Both humans and the LLM frequently reproduced the few-shot examples given in the annotation instructions, but diverged in their extensions: humans showed higher variability and occasionally omitted race altogether (4 prompts), whereas the LLM produced more systematic expansions that consistently included race.
Overall, the findings indicate that LLM-based bias detection consistently outperforms average human annotators and achieves coverage comparable to the consolidated union of multiple annotators. This suggests that LLMs can serve as reliable bias detectors, offering both scalability and consistency, and their use is justified as a practical alternative to extensive human annotation.
5 Experiments: Societal Bias Mitigation
In this section, we evaluate FairT2I against existing methods using automatic metrics and human studies, demonstrating the detection of diverse textual biases and the corresponding mitigation in outputs.
5.1 Automatic Evaluation
Evaluation dataset
To evaluate societal bias in professions, we use the same Stable Bias profession dataset (Luccioni_2023) introduced in section 4, which contains 131 occupations from the U.S. Bureau of Labor Statistics (BLS). We consider two target distributions: (1) the uniform distribution over categories defined in FairFace (karkkainen2021fairface), and (2) BLS occupational statistics. For the uniform distribution, we used the same subset of 50 randomly sampled occupations from this dataset as in section 4. The selected subset of prompts is listed in subsection B.1. For the BLS statistics, we use the five occupation prompts mentioned in naik2023social: CEO, doctor, computer programmer, nurse, and house keeper. As in zhang2023iti, each profession is prefixed with “A headshot of ” to form an input prompt.
Evaluation metrics
Following hirota2024saner; chuang2023debiasing, we measure bias in the generated outputs by computing the statistical parity (choi2020fair) between the empirical distribution and the ideal target distribution. In line with naik2023social, we detect faces in the generated images using dlib (king2015max) and then predict the attributes of each image using classifiers trained on the FairFace dataset (karkkainen2021fairface). Statistical parity is defined as the $\ell_{2}$‑distance between two probability vectors. Specifically, let $\mathbf{p}=(p_{1},p_{2},\dots,p_{n}),\quad\mathbf{q}=(q_{1},q_{2},\dots,q_{n})$ be two distributions over $n$ outcomes. Then the statistical parity between $\mathbf{p}$ and $\mathbf{q}$ is
$\mathrm{SP}(\mathbf{p},\mathbf{q})\;=\;\|\mathbf{p}-\mathbf{q}\|_{2}\;=\;\sqrt{\sum_{i=1}^{n}\bigl(p_{i}-q_{i}\bigr)^{2}}\,.$
A lower statistical parity indicates that the distribution of generated images more closely matches the target (fair) distribution.
For each prompt in our subset, we generate $200$ images, classify each by its FairFace attribute, and compute statistical parity. Further details appear in subsection C.1.
Methods for comparison
We compare FairT2I against existing debiasing methods that do not require any model finetuning:
  • Original: the unmodified Stable Diffusion model (sd), with no bias mitigation.
  • ENTIGEN (bansal2022well): appending an ethical injection phrase to the end of the prompt.
  • Fair Diffusion (friedrich2023fair): randomly sampling attributes from a predefined bias set and applying semantic guidance (brack2023sega).
Detailed hyperparameter settings for each method are provided in subsection C.2 and subsection C.3, respectively. To ensure a fair comparison, the T2I model was standardized to Stable Diffusion 1.5 (rombach2022high) across all baseline and proposed methods.
Implementation details
We used Stable Diffusion 1.5 (rombach2022high) as our T2I model. This choice was made because the hyperparameter settings and text encoders of the comparison methods were not compatible with the newer T2I architectures (sd3). We employed Claude 3.7-sonnet (anthropic2025claude37sonnet) for LLM-assisted bias detection, and GPT4o-mini (gpt4) to fuse the textual input $\bf{y}$ with the sensitive attribute $\bf{z}$. The parameter settings for image generation and the exact prompts supplied to the LLM are detailed in subsection C.4.
Table 2: Biased attributes associated with the input text “a headshot of a psychologist” as generated by LLM.
Attribute Category Associated Biases
Gender male, female, non-binary
Race White, Black, Asian, Hispanic, Middle Eastern, Other
Age young adult, middle-aged, elderly
Attire formal business attire, casual professional, white coat, sweater with blazer
Accessories glasses, no glasses, jewelry, no jewelry
Facial Hair beard, mustache, clean-shaven
Hair Style short, long, balding, styled, natural
Setting office, bookshelf background, plain background, clinical setting
Expression serious, smiling, attentive, compassionate
Targeting uniform distribution
Table 3 reports the SP scores comparing empirical and uniform distributions of gender and race in the Stable Bias Profession subset. FairT2I most closely matches the uniform baseline on both metrics. It demonstrates that, compared to existing methods, FairT2I generates data for each attribute with greater uniformity. See Appendix C for bar-chart comparisons and more details.
Table 2 summarizes the attribute categories (race, gender, age, background, pose, and more) and the biases detected by the LLM for the prompt “a headshot of a psychologist”. Figure 8 then presents randomly selected outputs from each method. While the baseline methods reproduce skewed or homogeneous attribute distributions, FairT2I addresses the biases listed in Table 2, yielding high-quality headshots with balanced diversity across race, gender, accessories (e.g., glasses), and backgrounds.
Targeting BLS statistics. Table 4 summarizes the gender and race SP scores between BLS statistics for CEO, computer programmer, doctor, nurse, and housekeeper prompts across four methods. FairT2I consistently achieves the lowest gender SP, outperforming ENTIGEN and FairDiffusion by large margins. For race SP, FairT2I yields the best scores on CEO, computer programmer, and housekeeper. These results confirm that FairT2I can also steer the generated distributions toward specified target distributions. See subsection C.8 for bar-chart comparisons and further details.
Table 3: Statistical Parity of Gender and Race Distributions by the original Stable Diffusion, ENTIGEN, FairDiffusion, and FairT2I (Ours).
Method Gender SP $\downarrow$ Race SP $\downarrow$
Original 0.2056 0.3907
ENTIGEN 0.0661 0.2001
FairDiffusion 0.2083 0.3137
FairT2I 0.0123 0.1864
Table 4: Gender Statistical Parity (top) and Race Statistical Parity (bottom) scores (lower is better) between BLS statistics for CEO, computer programmer, doctor, nurse and housekeeper in images from the original Stable Diffusion, ENTIGEN, FairDiffusion and FairT2I (Ours). For each metric and occupation, the best score is shown in bold.
Gender SP\ $\downarrow$
Method CEO comp. prog. doctor nurse housekeeper
Original 0.3637 0.1651 0.3843 0.1796 0.1459
ENTIGEN (bansal2022well) 0.3748 0.0569 0.1229 0.1210 0.0121
FairDiffusion (friedrich2023fair) 0.1956 0.1089 0.1138 0.1653 0.1597
FairT2I (Ours) 0.0710 0.0170 0.0550 0.0537 0.0121
Race SP\ $\downarrow$
Method CEO comp. prog. doctor nurse housekeeper
Original 0.6050 0.2213 0.2110 0.1971 1.0207
ENTIGEN (bansal2022well) 0.5004 0.6115 0.4727 0.4805 0.9341
FairDiffusion (friedrich2023fair) 0.1664 0.1484 0.0750 0.1285 0.3074
FairT2I (Ours) 0.0677 0.0885 0.1137 0.1541 0.2001
6 Experiments: Diversity Control
In this section, we evaluate FairT2I not only for its ability to mitigate societal bias but also for its capacity to enhance the diversity of generated images. We conduct both quantitative (automatic) and qualitative (human) evaluations to assess the effectiveness of our method.
6.1 Experimental Settings
Evaluation dataset
We employ the Parti Prompt dataset (yu2022scaling), which comprises over 1,600 diverse English prompts intended to rigorously evaluate T2I generation models and probe their limitations. In this work, we focus on the uniform prompt distribution, as no particular target distribution is assumed. For practical efficiency, we randomly sample 50 prompts from the full dataset as in section 4. The selected subset of prompts is listed in subsection B.1.
Models and baselines
We compare FairT2I against classifier-free guidance (CFG) (ho2022classifier) at three guidance scales (7.0, 4.0, 1.0), which modulate the trade‑off between image fidelity and diversity.
Implementation details
We use Stable Diffusion 3.5-large (sd3) as the T2I model. We employ Claude 3.7-sonnet (anthropic2025claude37sonnet) for LLM-assisted bias detection, and GPT-4o-mini (gpt4) to fuse textual input $\mathbf{y}$ with sensitive attribute $\mathbf{z}$. Generation hyperparameters and LLM prompts are in subsection D.2.
6.2 Automatic Evaluation
Automatic metrics and generation protocol
To assess image fidelity, we compute the Fréchet Inception Distance (FID) (heusel2017gans) on the COCO Karpathy test split (lin2014microsoft). To evaluate text–image alignment, we report CLIPScore (hessel2021clipscore) between each generated image and its input prompt. To quantify diversity, we embed all generated images into CLIP (clip) and BLIP2 (li2023blip) feature spaces, compute the covariance of embeddings, and use its trace as the diversity metric. For each prompt, we generate 200 images and compute these metrics accordingly.
Results
Table 5 presents quantitative evaluations in terms of FID, CLIPScore, CLIP Trace and BLIP2 Trace, illustrating the trade-off between image fidelity, text alignment and diversity. Under CFG 7.0 and CFG 4.0, FID values are low but Trace metrics remain small. By contrast, CFG 1.0 yields a high Trace value at the expense of elevated FID. Unlike these existing methods, FairT2I achieves low FID and high Trace without substantially reducing CLIPScore. Table 6 lists the attribute categories and biases detected by LLM for the prompt “an airplane flying into a cloud that looks like a monster.” in the Parti Prompts dataset. Figure 4 then presents the images generated by each method for the same prompt. While classifier-free guidance cannot achieve both image quality and diversity, FairT2I succeeds on both measures.
Table 5: Comparison of classifier-free guidance (CFG) at guidance scales 7.0, 4.0, and 1.0 and of FairT2I on the FID, CLIPScore, CLIP Trace, and BLIP2 Trace metrics. A guidance scale of 1.0 corresponds to generation without CFG.
Method FID\ $\downarrow$ CLIPScore\ $\uparrow$ CLIP Trace\ $\uparrow$ BLIP2 Trace\ $\uparrow$
CFG‑7.0 27.13 29.84 11.66 247.01
CFG‑4.0 26.46 30.09 11.87 250.24
No CFG (CFG‑1.0) 34.47 28.06 21.05 473.64
FairT2I (Ours) 26.24 28.35 19.18 454.49
Table 6: Attribute categories and their associated options generated by LLM for the input text “an airplane flying into a cloud that looks like monster.”.
Attribute Category Associated Biases
Type of Airplane commercial airliner, private jet, military aircraft, propeller plane, biplane
Cloud Appearance dark and menacing, fluffy and cute, wispy and ethereal
Monster Type dragon-like, humanoid, tentacled creature, multi-eyed beast, fanged monster
Time of Day daytime, sunset, night, dawn
Weather Conditions clear sky, stormy, partly cloudy
Viewing Angle from below, from above, side view, frontal view
Artistic Style photorealistic, cartoonish, dramatic, whimsical, ominous
Color Palette bright and colorful, dark and moody, pastel, monochromatic
Sky Background clear blue, sunset colors, stormy gray, gradient
Figure 4: Generated images for the input text “an airplane flying into a cloud that looks like monster.” by classifier-free guidance (CFG) at guidance scales 7.0, 4.0, and 1.0 and FairT2I (Ours). A guidance scale of 1.0 corresponds to generation without CFG.
6.3 Human Evaluation
To complement the automatic metrics, we conducted a human evaluation to assess the perceptual diversity, image quality, and text–image alignment of FairT2I in comparison with baseline methods.
Setup.
For each prompt in the 50-prompt subset of the Parti Prompt dataset, we generated nine images using each of four methods: CFG-1.0, CFG-4.0, CFG-7.0, and FairT2I. All images were arranged in $3\times 3$ grids per method, enabling side-by-side comparison across methods.
Task design
To keep the annotation time manageable, the 50 prompts were divided into five tasks, each containing ten prompts and the corresponding generations from all four methods. Each task required roughly 20 minutes to complete. We recruited a total of 100 annotators, resulting in 20 independent raters per prompt.
Procedure
Annotators were asked to rate every image on a five-point absolute Likert scale $(1=\text{worst},\,5=\text{best})$ along three criteria:
  1. 1.
    Diversity: How different the multiple images from the same model look for the same prompt. High diversity means the images show clearly different variations (e.g., viewpoint, layout, details), while low diversity means the images look nearly identical.
  2. 2.
    Image Quality: How clear, coherent, and visually convincing the image looks within its intended style. High-quality images are sharp, well-structured, and free from obvious AI artifacts (e.g., severe distortions, broken anatomy, unnatural textures), even if the content is fantastical or stylized (e.g., anime, surreal objects).
  3. 3.
    Text Alignment: How well the generated image matches the prompt in terms of objects, attributes, and relationships.
Detailed screenshots of the annotation interface, instructions, and answer forms are provided in Appendix D.5.
Results
Table 8 and Table 8 summarize the outcomes of the human evaluation. Inter-annotator agreement was modest but acceptable for aggregation (Krippendorff’s $\alpha$: Diversity $0.241$, Quality $0.141$, Alignment $0.136$). FairT2I achieved the highest mean ratings in both diversity and quality among all compared methods. For diversity, FairT2I obtained a mean score of $3.64\pm 1.10$, which was substantially higher than CFG-1.0 ($3.42$), CFG-4.0 ($2.70$), and CFG-7.0 ($2.74$). All corresponding one-sided Mann–Whitney U tests indicated statistically significant improvements ($p<0.001$), confirming that FairT2I produced perceptibly more diverse outputs than any baseline. For image quality, FairT2I also slightly outperformed all CFG variants ($3.83\pm 0.99$ vs. 2.782.783.753.75); the improvements over CFG-1.0 and CFG-4.0 were significant ($p<0.001$ and $p=4.7\times 10^{-5}$, respectively), while the difference from CFG-7.0 was not significant ($p=0.039$). In text–image alignment, FairT2I maintained competitive performance ($3.68\pm 1.12$), outperforming CFG-1.0 significantly ($p<0.001$) but not CFG-4.0 or CFG-7.0 ($p>0.9$). Overall, these results demonstrate that FairT2I improves perceived diversity and quality without compromising alignment, achieving a favorable balance between fidelity and variability that is consistent with the automatic evaluation trends.
Table 7: Human evaluation summary statistics for each task and model.
Criteria Model Mean Std.
Diversity CFG-1.0 3.416 1.120
CFG-4.0 2.697 1.154
CFG-7.0 2.735 1.226
FairT2I 3.643 1.105
Quality CFG-1.0 2.775 1.152
CFG-4.0 3.686 0.934
CFG-7.0 3.751 1.022
FairT2I 3.827 0.991
Alignment CFG-1.0 3.152 1.214
CFG-4.0 3.833 1.051
CFG-7.0 3.862 1.096
FairT2I 3.682 1.123
Table 8: One-sided Mann–Whitney U test results ($H_{1}$: FairT2I >> Baseline). Findings indicate whether FairT2I achieved a significantly higher score ($p<0.001$).
Criteria Baseline p-value Findings
Diversity CFG-1.0 $3.54\times 10^{-6}$ Significant $\uparrow$
CFG-4.0 $5.90\times 10^{-68}$ Significant $\uparrow$
CFG-7.0 $6.66\times 10^{-60}$ Significant $\uparrow$
Quality CFG-1.0 $1.44\times 10^{-87}$ Significant $\uparrow$
CFG-4.0 $4.74\times 10^{-5}$ Significant $\uparrow$
CFG-7.0 $3.93\times 10^{-2}$ Not significant
Alignment CFG-1.0 $4.97\times 10^{-23}$ Significant $\uparrow$
CFG-4.0 $9.98\times 10^{-1}$ Not significant
CFG-7.0 $1.00\times 10^{0}$ Not significant
7 FairT2I User Interface
Figure 5: Our web application for bias-aware image generation. (1) Users first input a text prompt for image generation. (2) The LLM detects potential societal biases or implicit characteristics and presents them in a tabular format. (3) Users can edit the table to adjust categories or attributes as desired. (4) The T2I model then generates images based on the edited table, enabling debiased image generation. Please also refer to the supplemental video.
To demonstrate the feasibility of our FairT2I framework, we developed an interactive web interface that enables users to experience bias-aware image generation (see Figure 5 and the supplemental video). The interface is implemented using Gradio [^2] https://www.gradio.app/. This interface allows users not only to identify and mitigate potential societal biases but also to promote and control diversity in the generated outputs.
7.1 Input and Bias Detection
Users begin by entering a text prompt (e.g., “A portrait photo of a computer programmer”). Upon submission, the LLM analyzes the prompt and detects potential bias-relevant categories such as gender, race, age, appearance, and setting. The detected categories and their representative attributes (e.g., male/female/non-binary for gender, White/Asian/Black for race) are automatically populated in a tabular format. This step exposes the model’s interpretive assumptions and encourages users to reflect on implicit biases present in seemingly neutral prompts.
7.2 Interactive Bias Editing (Optional)
The detected table is fully editable, enabling users to modify the bias representation before image generation. Users can (1) add or remove entire categories (for example, delete posture or add occupation); (2) introduce new attributes within a category (for example, add “teen” under age); and (3) adjust attribute weights using sliders to control the sampling ratio of each attribute during prompt generation. By default, attribute weights are initialized to a uniform distribution, which ensures equal sampling probability across all detected attributes. Users can also freely modify these weights, for instance to match real world statistics, specific target distributions, or any other desired sampling strategy, thereby providing fine grained control over the generative bias.
7.3 Image Generation and Visualization
After making adjustments, users press the Generate Images button to produce multiple debiased outputs. The system constructs a set of balanced prompts based on the edited table and inputs them into the underlying T2I model. The generated images are then displayed in a grid layout, each accompanied by its rewritten prompt text. This visualization helps users see how modifying textual bias structures influences the resulting visual outputs.
8 Limitations
LLM bottlenecks
As shown in Figure 9, FairT2I generates fewer images for Southeast Asian and Indian categories compared to other classes. This is partly because the LLM does not detect Southeast Asian or Indian as race categories. Notably, in the user study reported in section 4, only 3 out of 25 participants included Indian in the race category, and none mentioned Southeast Asian. This suggests that such categories are underrepresented not only in LLM outputs but also in human conceptualization, which can further reinforce detection biases. By explicitly supplying these attributes as few-shot examples in the LLM prompt, this issue can be mitigated.
Text encoder bottlenecks
FairT2I reflects latent attributes in image generation via explicit prompt modifications. As demonstrated by hirota2024saner, information such as gender that is explicitly specified is accurately captured by the text encoder in the generated images. However, long prompts and complex logical relations (tang2023lemons; zhang2023iti) may not be reflected in the outputs. This limitation could potentially be addressed by using a long-context model (zhang2024long) or a vision-language model-based text encoder (chen2025multimodal).
Understanding users and interactions
While our interface demonstrates the feasibility of bias-aware T2I generation, it remains an early prototype in terms of human–AI interaction design. At present, the system allows users to view and edit detected bias categories and attributes; however, the way users interpret these biases and decide to modify them remains unclear. Future work could involve conducting dedicated interaction studies to investigate how users engage with detected potential biases, thereby obtaining general insights for designing bias-aware T2I generation interactions. This includes understanding how users interpret and manipulate detected biases, as well as the strategies they adopt to balance fairness and creative intent.
9 Conclusions
In this paper, we have presented FairT2I, a novel approach to debiasing T2I models by decomposing score functions using latent variable guidance and leveraging LLMs for bias detection and attribute resampling. We confirmed FairT2I can effectively debias T2I models without the need for model fine-tuning or tedious hyperparameter tuning. Furthermore, FairT2I can be used not only for societal bias mitigation but also to improve and control diversity.
Ethical Considerations
All authors carefully reviewed and followed the 2021 ACM Publications Policy on Research Involving Human Participants and Subjects and the ACM Code of Ethics, and the study adhered to the principles of respect, fairness, and risk minimization.
Acknowledgments
We are grateful to Yusuke Hirota for his insightful discussions, guidance on the experiments, and careful review of the manuscript.
Appendix A Proof of Proposition 1
Proposition 1.
Let $\mathbf{y}$ represent the input text, $\mathbf{x}$ the generated image, and $\mathbf{z}$ a discrete latent variable taking values in a finite set $\mathcal{Z}$, and assume conditional independence between $\mathbf{x}$ and $\mathbf{z}$ given $\mathbf{y}$, the following equation holds:
$\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{y})=\sum_{z\in\mathcal{Z}}p(\mathbf{z}=z\mid\mathbf{y})\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y}).$ (2)
Proof.
We start from the left-hand side (LHS) of the equation stated in Proposition 1.
$\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{y})$ $=\frac{1}{p(\mathbf{x}\mid\mathbf{y})}\nabla_{\mathbf{x}}p(\mathbf{x}\mid\mathbf{y})$ (Using the chain rule for logarithms)
$=\frac{1}{p(\mathbf{x}\mid\mathbf{y})}\nabla_{\mathbf{x}}\sum_{z\in\mathcal{Z}}p(\mathbf{x},\mathbf{z}=z\mid\mathbf{y})$ $\text{(Marginalizing over }\mathbf{z}\text{)}$
$=\frac{1}{p(\mathbf{x}\mid\mathbf{y})}\nabla_{\mathbf{x}}\sum_{z\in\mathcal{Z}}p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y})p(\mathbf{z}=z\mid\mathbf{y})$ (Using the definition of conditional probability)
$=\frac{1}{p(\mathbf{x}\mid\mathbf{y})}\sum_{z\in\mathcal{Z}}p(\mathbf{z}=z\mid\mathbf{y})\nabla_{\mathbf{x}}p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y})$ $\text{(Linearity of }\nabla_{\mathbf{x}}\text{)}$
$=\sum_{z\in\mathcal{Z}}\frac{p(\mathbf{z}=z\mid\mathbf{y})}{p(\mathbf{x}\mid\mathbf{y})}\nabla_{\mathbf{x}}p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y}).$
Then we apply Bayes’ theorem,
$p(\mathbf{z}=z\mid\mathbf{x},\mathbf{y})=\frac{p(\mathbf{z}=z\mid\mathbf{y})p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y})}{p(\mathbf{x}\mid\mathbf{y})},$
to obtain
$\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{y})$ $=\sum_{z\in\mathcal{Z}}\frac{p(\mathbf{z}=z\mid\mathbf{x},\mathbf{y})}{p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y})}\nabla_{\mathbf{x}}p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y})$
$=\sum_{z\in\mathcal{Z}}p(\mathbf{z}=z\mid\mathbf{x},\mathbf{y})\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y}).$ (Using the chain rule for logarithms)
Finally, we apply the conditional independence assumption given in the proposition, $p(\mathbf{z}=z\mid\mathbf{x},\mathbf{y})=p(\mathbf{z}=z\mid\mathbf{y})$:
$\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{y})=\sum_{z\in\mathcal{Z}}p(\mathbf{z}=z\mid\mathbf{y})\nabla_{\mathbf{x}}\log p(\mathbf{x}\mid\mathbf{z}=z,\mathbf{y}).$
Appendix B Details for Bias Detection Experiment
B.1 Evaluation Dataset Details
For efficient evaluation, we used a subset of 50 randomly selected prompts from each dataset: profession names from the Stable Bias dataset [Luccioni_2023] and textual prompts from the Parti Prompt dataset [yu2022scaling]. The selected subsets are listed in Table 9 and Table 11, respectively.
Table 9: A 50 profession subset of the stable bias professions dataset used in the experiment. The subset is randomly selected from the full dataset.
Professions
aerospace engineer aide author bartender
carpenter cashier CEO civil engineer
cleaner coach compliance officer cook
dental assistant dentist detective electrical engineer
engineer facilities manager fast food worker file clerk
graphic designer hairdresser head cook health technician
industrial engineer interior designer interviewer inventory clerk
jailer machinery mechanic manicurist massage therapist
medical records specialist mental health counselor metal worker office clerk
painter payroll clerk physical therapist plane mechanic
postal worker psychologist purchasing agent repair worker
roofer sales manager sheet metal worker social worker
underwriter welder
Table 10: Gender and race distributions from BLS 2024 statistics. Race percentages have been normalized so that they sum to 100 %.
Occupation Female (%) Male (%) White (%) Black (%) Asian (%) Hispanic (%)
CEO 33.0 67.0 82.2 5.8 6.1 5.8
Doctor 44.5 55.5 64.6 7.0 22.2 6.2
Computer Programmer 17.8 82.2 65.7 8.3 16.0 10.1
Nurse 86.8 13.2 67.0 14.7 9.1 9.1
Housekeeper 87.7 12.3 51.3 10.2 3.1 35.3
Table 11: A 50 prompt subset of the Parti Prompt dataset used in the experiment. The subset is randomly selected from the full dataset.
No. Description No. Description
1. a pile of cash on a wooden table 26. a paranoid android freaking out and jumping into the air because it is surrounded by colorful Easter eggs
2. five frosted glass bottles 27. a musical note
3. a view of the Big Dipper in the night sky 28. Anubis wearing sunglasses and sitting astride a hog motorcycle
4. a cow 29. a photograph of a squirrel holding an arrow above its head and holding a longbow in its left hand
5. a Ferrari Testarossa in front of the Kremlin 30. an elder politician giving a campaign speech
6. a rowboat 31. a hot air balloon with a yin-yang symbol
7. a portrait of a postal worker who has forgotten their mailbag 32. overhead view of three people looking down at the street from the top of a tall building
8. a helicopter hovering over Times Square 33. meaning of life
9. a t-shirt with Carpe Diem written on it 34. a laptop no letters on its keyboard
10. a canal in Venice 35. a glass of orange juice with an orange peel stuck on the rim
11. trees seen through a car window on a rainy day 36. a yellow tiger with blue stripes
12. a grumpy porcupine handing a check for \$10 37. cash
13. a watermelon chair 38. matching socks with cute cats on them
14. a triangle with a smiling face 39. five chairs
15. Four deer surrounding a moose. 40. an airplane flying into a cloud that looks like monster
16. an old red truck parked by the geyser Old Faithful 41. a long-island ice tea cocktail next to a napkin
17. a tree surrounded by flowers 42. a white robot passing a soccer ball to a red robot
18. A photo of a Ming Dynasty vase on a leather topped table. 43. a black towel
19. a chair 44. a capybara
20. a wooden post 45. artificial intelligence
21. Jupiter rises on the horizon. 46. Portrait of a tiger wearing a train conductor’s hat and holding a skateboard that has a yin-yang symbol on it. charcoal sketch
22. A raccoon wearing formal clothes 47. Anime illustration of the Great Pyramid sitting next to the Parthenon under a blue night sky of roiling energy
23. A bowl of soup that looks like a monster with tofu says deep learning 48. a chess queen to the right of a chess knight
24. the Mona Lisa in the style of Minecraft 49. a Tyrannosaurus Rex roaring in front of a palm tree
25. a sphere 50. Portrait of a gecko wearing a train conductor’s hat and holding a flag that has a yin-yang symbol on it. Child’s crayon drawing.
B.2 Pilot Study Details
We conducted a pilot comparison of four LLMs: GPT-4o gpt4, Claude-3.7-Sonnet anthropic2025claude37sonnet, LLaMA-3.3 grattafiori2024llama3herdmodels, and DeepSeek-V3 liu2024deepseek, to assess their ability to detect potential bias-relevant categories and attributes in text prompts. The comparison was performed using two evaluation subsets: 50 prompts sampled from the Stable Bias dataset and another 50 prompts from the Parti Prompt dataset, as listed in subsection B.1.
For each model, we calculated the average number of detected categories per prompt and the average number of attributes per category. As summarized in Figure 6, Claude3.7-Sonnet consistently detected a broader range of bias-relevant categories and attributes compared to the other LLMs, followed by LLaMA-3.3 and DeepSeek-V3, while GPT-4o showed the most conservative detection behavior.
Quantitatively, on the Stable Bias subset, Claude detected $8.46\pm 1.49$ categories with $5.54\pm 0.92$ attributes per category, outperforming GPT-4o ($5.06\pm 1.04$ and $3.51\pm 0.80$, respectively), LLaMA 3.3 ($7.42\pm 1.37$ and $4.25\pm 0.56$), and DeepSeek-V3 ($6.82\pm 1.61$ and $4.77\pm 0.70$). Similarly, on the Parti Prompt subset, Claude achieved $9.06\pm 0.81$ categories and $4.54\pm 0.46$ attributes per category, again exceeding GPT-4o ($5.34\pm 0.59$, $3.32\pm 0.28$), LLaMA-3.3 ($7.78\pm 0.58$, $3.88\pm 0.38$), and DeepSeek-V3 ($7.00\pm 1.04$, $4.07\pm 0.34$).
Table 12 further illustrates qualitative differences among the models using the prompt “A CEO” as an example. Claude 3.7 Sonnet identifies a richer set of categories, such as body type, accessories, and facial expression, which were often omitted by other models. This broader detection suggests that Claude may capture subtler implicit biases embedded in occupational or descriptive text, offering a more comprehensive representation of latent societal stereotypes.
Figure 6: The number of detected categories and attributes per category for Stable Bias (left) and Parti Prompt (right) datasets. Bars show Claude3.7-sonnet (light blue), GPT-4o (orange), LlaMA-3.3 (green), and DeepSeek-V3.2 (blue) with error bars indicating one standard deviation.
Table 12: Detected bias categories and corresponding attributes by different LLMs for the input prompt “A CEO”.
Category Claude 3.7 Sonnet GPT-4o LLaMA 3.3 DeepSeek V3.2
Gender male, female, non-binary male, female, non-binary male, female, non-binary male, female, non-binary
Age young, middle-aged, elderly young adult, middle-aged, elderly young adult, middle-aged, elderly middle-aged, elderly, young adult
Race White, Black, Asian, Hispanic, Middle Eastern, Indigenous White, Asian, Black, Hispanic White, Asian, Black, Hispanic, Middle Eastern White, Asian, Black, Hispanic, Middle Eastern
Appearance / Attire formal attire, business suit, casual business attire, professionally dressed business suit, casual wear, formal wear formal, business casual, luxurious, conservative, modern well-dressed, formal attire, casual attire, professional
Body type / Physical ability slim, athletic, average, plus-sized able-bodied, uses mobility aid
Setting office, boardroom, corporate headquarters, tech company, factory floor boardroom, office, conference room corporate, industrial, academic
Background urban, rural ivy league educated, self-made, inherited position, career climber
Pose sitting at desk, standing, presenting, in meeting, speaking
Accessories glasses, watch, briefcase, laptop, smartphone, tablet
Facial expression / Personality serious, smiling, confident, authoritative, friendly confident, serious, smiling, determined, approachable authoritative, charismatic, decisive, collaborative
Industry / Environment tech, finance, healthcare, manufacturing, services technology, finance, healthcare, manufacturing, retail
Nationality American, European, Asian, Global
B.3 User Study Details
We conducted a user study using the crowdsourcing platform Prolific [^3] https://www.prolific.com/ to collect annotations from diverse participants. Prolific allows researchers to recruit participants according to specific demographic quotas. In our case, we applied sex and ethnicity quotas to ensure balanced participation and obtain responses from a more diverse pool of annotators.
All responses were collected using a Google Form [^4] https://docs.google.com/forms/. Given a set of text prompts, participants were instructed to list potential biases or implicit characteristics in the form of categories and attributes. To facilitate understanding, the instruction included several example answers illustrating how participants should format their responses, so that they could easily grasp our intended task. The task interface shown to participants is illustrated in Figure 7. Furthermore, we enforced a regular-expression-based answer format to ensure consistent and well-structured annotations.
Figure 7: The task interface shown to participants, including the task description, sample prompts, and entry fields for submitting responses.
We prepared a total of 100 text prompts, consisting of 50 prompts from the Stable Bias dataset [Luccioni_2023] and 50 prompts from the PartiPrompts dataset [yu2022scaling], as detailed in subsection B.1. To maintain reasonable annotation time, these 100 prompts were divided into five separate tasks, each containing 20 prompts. Different tasks were assigned to different participants to avoid repetition.
For each task, we recruited five independent annotators, resulting in a total of 25 participants. This configuration kept the estimated completion time around 30 minutes per participant. Each participant received a compensation of £4.5, corresponding to an hourly rate of approximately £9.0, which aligns with fair payment standards on Prolific.
Appendix C Details for Societal Bias Mitigation Experiment
C.1 Evaluation Metric Details
The FairFace classifier outputs two gender categories, male and female, and seven race categories, White, Southeast Asian, Middle Eastern, Latino_Hispanic, Indian, East Asian, and Black. Based on these classification results, we compute the statistical parity against the target distributions. Table 10 shows the 2024 BLS employment statistics [^5] https://www.bls.gov/cps/cpsaat11.htm for CEO, computer programmer, doctor, nurse, and housekeeper. The original statistics report the percent of total employed for women, White, Black or African American, Asian, and Hispanic or Latino. To construct the BLS target distribution, we assigned the proportion of women to the probability of female and one minus that proportion to male in the gender category. For the race categories, we normalized the proportions for White, Black or African American, Asian, and Hispanic or Latino so that they sum to one [^6] NOTE: Estimates for the above race groups (White, Black or African American, and Asian) do not sum to totals because data are not presented for all races. Persons whose ethnicity is identified as Hispanic or Latino may be of any race. and used this as the target distribution. Because the FairFace classifier does not include a generic Asian category, we combined the probabilities of Southeast Asian and East Asian to represent Asian when computing statistical parity.
C.2 ENTIGEN Setup Details
As an ethical intervention to promote equitable judgment in T2I models, we added “, if all individuals can be ¡PROFESSION¿ irrespective of their gender and race” to the end of the input prompt.
C.3 Fair Diffusion Setup Details
For gender editing, we predefined the prompts [“male person”, “female person”, “non-binary person”], and for race editing, the prompts [“white person”, “black person”, “asian person”, “latino person”, “indian person”]. During sampling, one gender prompt and one race prompt were drawn according to the target distribution, and reverse_editing_direction=True was used to enhance the selected attribute while suppressing the others. All editing operations employed the following hyperparameters (as in the original implementation):
  • edit_warmup_steps = 1
  • edit_guidance_scale = 3.0
  • edit_threshold = 0.95
  • edit_weights = 1.0
  • edit_momentum_scale = 0.3
  • edit_momentum_beta = 0.6
Except for these settings, all other parameters matched those of the original work friedrich2023fair.
C.4 FairT2I Setup Details
T2I model sampling was performed using PNDMScheduler with 50 sampling steps. The guidance_scale for classifier-free guidance was set to 7.5. During inference, we set the batch size to 200 and image resolution 512 $\times$ 512, and a single inference on one NVIDIA H100 GPU took approximately 4 minutes.
For bias detection, we used claude-3-7-sonnet-20250219, and for prompt rewriting, we used gpt-4o-mini-2024-07-18.
The prompt used for LLM-assisted bias detection is as follows:
Please read the following task description carefully. If you do not follow the specified format, your answer will not be accepted.
Task
Read the descriptive text carefully. Think about what kinds of biases that could appear in an image (stereotypes, assumptions, or implicit characteristics) may be implied.
Important Rules
You must create at least 2 categories.
Each category must list 2 or more attributes.
Try to list as many categories and attributes as you can think of.
Format must follow:
 ”category”: [”attribute1”, ”attribute2”, …],
 ”category”: [”attribute1”, ”attribute2”, …],
Do not include any text outside of the JSON output.
If you do not follow this format, your answer will be rejected.
Examples
If the text is: “a firefighter”
Correct format:
 ”gender”: [”male”, ”female”, ”non-binary”],
 ”age”: [”young adult”, ”middle-aged”, ”elderly”],
 ”race”: [”White”, ”Asian”, ”Black”, ”Hispanic”],
Incorrect format:
 ”race”: ”Asian”
If the text is: “a quiet forest”
Correct format:
 ”time_of_day”: [”morning”, ”afternoon”, ”evening”, ”night”, ”dawn”, ”dusk”],
 ”season”: [”spring”, ”summer”, ”autumn”, ”winter”],
 ”weather”: [”sunny”, ”cloudy”, ”rainy”, ”snowy”, ”foggy”, ”misty”],
Incorrect format: morning
The following prompt was used for the LLM-assisted fusion of the textual input $\mathbf{y}$ and the latent attribute $\mathbf{z}$:
Please read the following task description carefully. If you do not follow the specified format, your answer will not be accepted.
Task
Rewrite the original prompt so that it naturally integrates all specified attributes.
Important Rules
Output must contain only the rewritten prompt text. Do not add, expand, or explain beyond the original prompt and attributes. Do not include extra descriptive words, embellishments, or commentary. Do not include explanations, metadata, or formatting outside of the rewritten prompt. The rewritten sentence must be as concise as possible, while preserving the original meaning and smoothly integrating all attributes. Format
<rewritten prompt text>
If you do not follow this format, your answer will be rejected.
Examples
If the input is:
 Original prompt: ”A portrait of a person reading a book”
 Attributes to include: ”gender: female, age: elderly”
Correct format:
 A portrait of an elderly female person reading a book
Incorrect format:
 A detailed portrait depicting an elderly female person seated in a quiet and serene setting, holding a book gently in her hands while reading it with focused attention and calm concentration.
C.5 Statistical Significance Test Methodology
To determine if FairT2I achieved statistically significantly lower SP scores compared to another method (denoted as Method A), we employed a one-sided non-parametric bootstrap hypothesis test with $N_{boot}=1000$ iterations. In each iteration $i$, demographic classifications were resampled with replacement from the set of generated images for Method A and FairT2I, maintaining the original sample sizes from which the SP scores were computed. Specifically:
  • For the uniform distribution target (see subsection C.6), overall SP scores were initially calculated from demographic classifications of all $50\text{ occupations}\times 200\text{ images/occupation}=10,000$ images generated per method. The bootstrap procedure then resampled from these $10,000$ image classifications for each method to generate bootstrap SP scores.
  • For the BLS statistics target (see subsection C.8), SP scores were initially calculated for each occupation based on the $n=200$ images generated per occupation per method. The bootstrap procedure then resampled from the $200$ image classifications for a given occupation for each method.
SP scores were then calculated for these bootstrap samples ($\mathrm{SP}_{A,i}^{*}$ for Method A and $\mathrm{SP}_{B,i}^{*}$ for FairT2I, where B denotes FairT2I). The difference $D_{i}^{*}=\mathrm{SP}_{A,i}^{*}-\mathrm{SP}_{B,i}^{*}$ was computed for each iteration. The $p$-value was estimated as the proportion of these bootstrap differences $D_{i}^{*}$ that were less than or equal to zero: $p=\frac{\sum_{i=1}^{N_{boot}}\mathbb{I}(D_{i}^{*}\leq 0)}{N_{boot}}$, where $\mathbb{I}(\cdot)$ is the indicator function. This $p$-value tests the null hypothesis $H_{0}:\mathrm{SP}_{\text{FairT2I}}\geq\mathrm{SP}_{\text{Method A}}$ against the alternative hypothesis $H_{A}:\mathrm{SP}_{\text{FairT2I}}<\mathrm{SP}_{\text{Method A}}$. $p<0.05$ indicates that FairT2I is significantly better than Method A for the given SP metric.
C.6 Detailed Results of Uniform Distribution Target
The overall SP scores for gender and race when targeting a uniform distribution are presented in Table 13. The statistical significance of FairT2I’s scores compared to other methods was assessed using the bootstrap test described in subsection C.5.
Table 13: Gender SP (left) and Race SP (right) scores (lower is better) from the uniform distribution. For each metric, the best score is shown in bold. For Original, ENTIGEN, and FairDiffusion, the $p$-value in parentheses ($p$-val) indicates if FairT2I (Ours) is significantly better than that method ($p<0.05$ highlighted in bold for $\mathrm{SP}_{\text{FairT2I}}<\mathrm{SP}_{\text{Method}}$).
Method Gender SP $\downarrow$ Race SP $\downarrow$
Original 0.2056 (<<0.0001) 0.3907 (<<0.0001)
ENTIGEN 0.0661 (<<0.0001) 0.2001 (<<0.0001)
FairDiffusion 0.2083 (<<0.0001) 0.3137 (<<0.0001)
FairT2I (Ours) 0.0123 0.1864
Figure 8: Generated images for the input text “a headshot of a psychologist” by the original Stable Diffusion, ENTIGEN, FairDiffusion, and FairT2I (Ours).
C.7 Evaluation Results
Figure 9: Comparison of the target uniform distribution and empirical distributions produced by original Stable Diffusion, ENTIGEN, FairDiffusion and FairT2I (Ours) for gender (left) and race (right), shown as stacked bar charts.
Discussion of uniform distribution results
As shown in Table 13, when targeting a uniform demographic distribution across all occupations, FairT2I (Ours) achieved the best overall SP scores for both Gender SP (0.0123) and Race SP (0.1864). The statistical significance tests, detailed in subsection C.5, confirm these improvements. FairT2I is shown to be significantly better than the Original method (Gender SP: 0.2056, Race SP: 0.3907), ENTIGEN (Gender SP: 0.0661, Race SP: 0.2001), and FairDiffusion (Gender SP: 0.2083, Race SP: 0.3137) for both metrics, with all associated $p$-values being less than 0.000. This indicates a strong and statistically robust improvement in fairness for overall image generation when a uniform distribution is the goal.
Stacked Bar-Chart Visualization
Figure 9 presents a stacked bar-chart comparison between the target uniform distribution and the empirical distributions produced by Stable Diffusion, ENTIGEN, Fair Diffusion, and our proposed FairT2I. Among these methods, FairT2I demonstrates the closest alignment with the target distribution.
C.8 Detailed Results of BLS Statistics Target
For the BLS statistics target, FairT2I’s performance in achieving lower SP scores per occupation was also evaluated against other methods. The statistical significance of these comparisons was determined using the one-sided non-parametric bootstrap hypothesis test detailed in subsection C.5. The SP scores and the corresponding $p$-values from these tests are presented in Table 14.
Table 14: Gender SP (top) and Race SP (bottom) scores (lower is better) when targeting BLS statistics. For each metric and occupation, the best score is shown in bold. For Original, ENTIGEN, and FairDiffusion, the $p$-value in parentheses ($p$-val) indicates if FairT2I (Ours) is significantly better than that method ($p<0.05$ highlighted in bold for $\mathrm{SP}_{\text{FairT2I}}<\mathrm{SP}_{\text{Method}}$).
Gender SP\ $\downarrow$
Method CEO comp. prog. doctor nurse housekeeper
Original 0.3637 (0.000) 0.1651 (0.000) 0.3843 (0.000) 0.1796 (0.001) 0.1459 (0.000)
ENTIGEN 0.3748 (0.000) 0.0569 (0.254) 0.1229 (0.123) 0.1210 (0.044) 0.0121 (0.547)
FairDiffusion 0.1956 (0.022) 0.1089 (0.044) 0.1138 (0.206) 0.1653 (0.004) 0.1597 (0.000)
FairT2I (Ours) 0.0710 0.0170 0.0550 0.0537 0.0121
Race SP\ $\downarrow$
Method CEO comp. prog. doctor nurse housekeeper
Original 0.6050 (0.000) 0.2213 (0.003) 0.2110 (0.016) 0.1971 (0.234) 1.0207 (0.000)
ENTIGEN 0.5004 (0.000) 0.6115 (0.000) 0.4727 (0.000) 0.4805 (0.000) 0.9341 (0.000)
FairDiffusion 0.1664 (0.019) 0.1484 (0.027) 0.0750 (0.752) 0.1285 (0.714) 0.3074 (0.008)
FairT2I (Ours) 0.0677 0.0886 0.1137 0.1540 0.2001
Discussion of BLS statistics results
The analyses in Table 14 indicate that FairT2I generally achieves lower SP scores for both gender and race attributes compared to existing methods when targeting BLS statistics. For Gender SP, FairT2I consistently demonstrates statistically significant improvements ($p<0.05$) over the Original method across all five occupations. It also shows significant advantages over FairDiffusion in most occupations (CEO, computer programmer, nurse, housekeeper). Compared to ENTIGEN, FairT2I provides significant improvements in two occupations (CEO, nurse), while for others, including a tied best performance with ENTIGEN for housekeeper, the differences were not statistically significant. Regarding Race SP, FairT2I shows robust and statistically significant improvements over ENTIGEN across all occupations. Against the Original method, FairT2I is significantly better in four out of five occupations. When compared with FairDiffusion, FairT2I offers significant gains for three occupations (CEO, computer programmer, housekeeper). Notably, FairDiffusion achieved lower Race SP scores for the doctor and nurse occupations, where FairT2I did not show a significant advantage.
Stacked bar chart visualization
Figure 10 presents stacked bar charts comparing demographic distributions of target BLS statistics and empirical distributions across five occupations. Visually, FairT2I generally demonstrates a closer alignment to the target BLS distributions for both gender and race across most occupations.
Figure 10: Comparison of the target BLS statistics and empirical distributions produced by original Stable Diffusion, ENTIGEN, FairDiffusion and FairT2I (Ours) for gender (left) and race (right), shown as stacked bar charts.
Appendix D Details for Diversity Control Experiment
D.1 Evaluation Metrics Details
This section provides further details on the implementation of the evaluation metrics used in Section 6 to assess image fidelity, text-image alignment, and diversity. For all trace-based diversity metrics, we generated 200 images for each of the 50 sampled Parti Prompts. Image features were extracted from these 200 images, and the trace of their covariance matrix was computed.
FID.
We calculated FID heusel2017gans to assess image fidelity against the COCO Karpathy test split [lin2014microsoft]. We used the PyTorch implementation of FID from the pytorch-fid library [^7] https://github.com/mseitzer/pytorch-fid.
CLIPScore.
To evaluate text-image alignment, we computed CLIPScore hessel2021clipscore.
  • Model and Library: We utilized the OpenAI CLIP ViT-B/32 model, accessed via the official clip Python package.
  • Feature Extraction: For each (prompt, generated image) pair, the prompt text was encoded using model.encode_text() and the image was encoded using model.encode_image(). Both encoders produce L2-normalized features.
  • Score Calculation: The CLIPScore for a single pair is the dot product of their respective normalized feature embeddings.
  • Aggregation: The final reported CLIPScore for a given prompt is the average score across all 200 images generated for that prompt.
CLIP Trace.
To quantify diversity using CLIP clip features, we calculated the trace of the covariance matrix of image embeddings.
  • Model and Library: We used the OpenAI CLIP ViT-B/32 model, via the clip Python package.
  • Feature Extraction: For each of the 50 sampled prompts, we generated 200 images. All 200 images were encoded into L2-normalized feature vectors using model.encode_image() with a batch size of 32. This resulted in a set of 200 feature vectors per prompt.
  • Trace Calculation: For each prompt, we computed the covariance matrix of these 200 feature vectors using numpy.cov(features, rowvar=False). The CLIP Trace diversity metric is the trace of this covariance matrix, calculated using numpy.trace().
BLIP2 Trace.
Similar to CLIP Trace, we also quantified diversity using BLIP-2 li2023blip features.
  • Model and Library: We employed the Salesforce/blip2-opt-2.7b model, accessed via the Hugging Face transformers library, specifically using Blip2Processor for preprocessing and Blip2Model for feature extraction.
  • Feature Extraction: For each prompt, the 200 generated images were processed. Image features were obtained from the pooler_output of model.get_image_features().
  • Trace Calculation: Analogous to CLIP Trace, for each prompt, we computed the covariance matrix of the 200 BLIP-2 image feature vectors and then calculated its trace.
D.2 FairT2I Setup Details
T2I model sampling was performed using FlowMatchEulerDiscreteScheduler with 28 sampling steps. The guidance_scale for classifier-free guidance was set to 4.0. During inference, we set the batch size to 200 and image resolution 1024 $\times$ 1024, and a single inference on one NVIDIA H100 GPU took approximately 4 minutes. For bias detection, we used claude-3-7-sonnet-20250219, and for prompt rewriting, we used gpt-4o-mini-2024-07-18.
The same prompt in subsection C.4 was used for LLM-assisted bias detection is as follows. We utilized the same prompt for integration of the input text $\mathbf{y}$ and the sensitive attribute $\mathbf{z}$ with subsection C.4.
D.3 Statistical Significance Test Methodology
To assess the statistical significance of the differences observed in diversity (BLIP2 Trace, CLIP Trace) and text-image alignment (CLIP Score) metrics (see section 6 in the main paper), we performed one-sided Mann-Whitney U tests. This non-parametric test was chosen as it does not assume a specific distribution (e.g., normality) for the data and is suitable for comparing two independent samples of per-prompt scores. For each comparison between two methods (Group 1 and Group 2, as detailed in Table 15), we tested the alternative hypothesis ($H_{A}$) that the scores from Group 1 are stochastically greater than those from Group 2. The detailed results of these tests, including $p$-values and findings, are presented in Table 15. Note that statistical tests for FID scores were not performed due to the high computational cost associated with generating multiple full sets of images for each condition required for robust FID variance estimation. FID scores are reported as single values based on one comprehensive generation run per method.
D.4 Discussion of Diversity, Alignment, and Fidelity Results
The statistical analyses presented in Table 15, along with reported FID scores, provide insights into FairT2I’s performance in terms of image diversity, text-image alignment, and image fidelity, relative to baseline Stable Diffusion models with varying classifier-free guidance (CFG) scales.
Diversity (Trace Scores)
FairT2I (itself using a CFG scale of 4.0) demonstrates a statistically significant increase in diversity, as measured by both BLIP2 Trace and CLIP Trace, when compared to the original Stable Diffusion model using higher or equivalent CFG scales (i.e., CFG 7.0 and CFG 4.0). Specifically, for both trace metrics, FairT2I’s scores were significantly higher (all $p<0.00001$, except for BLIP2 Trace vs. Original CFG 4.0 where $p=0.000002$). When compared to the original model with a very low CFG scale (CFG 1.0), which typically yields high diversity, the results are nuanced. While the original CFG 1.0 model achieved numerically higher mean trace scores for both BLIP2 Trace and CLIP Trace compared to FairT2I, these differences were not found to be statistically significant (BLIP2 Trace $p=0.255$; CLIP Trace $p=0.069$ when testing $H_{A}$: Original CFG 1.0 >> FairT2I). This suggests FairT2I achieves a level of diversity that is statistically comparable to that of the highly diverse CFG 1.0 setting.
Text-image alignment (CLIP Score)
For CLIP Scores, the original model with CFG scales of 7.0 and 4.0 yielded statistically significantly higher scores than FairT2I ($p=0.021$ and $p=0.008$, respectively, when testing $H_{A}$: Original >> FairT2I). This indicates a slight, yet statistically significant, decrease in text-image alignment for FairT2I when compared to these higher CFG scale baselines. However, when comparing FairT2I to the original model with CFG 1.0, there was no statistically significant evidence that Original CFG 1.0’s CLIP Score was higher than FairT2I’s ($p=0.664$). Numerically, FairT2I’s mean CLIP Score was slightly higher than that of Original CFG 1.0. This suggests a trade-off: increasing diversity via FairT2I might involve a small compromise in text-image alignment relative to high-guidance settings, but the alignment remains competitive, especially when compared to other high-diversity configurations like CFG 1.0.
Image Fidelity (FID) and overall balance
While statistical tests were not performed for FID due to computational expense, the reported scores (Original CFG 7.0: 27.13, Original CFG 4.0: 26.46, Original CFG 1.0: 34.47, FairT2I: 26.24) indicate that FairT2I achieves the best image fidelity (lowest FID score). It surpasses or matches the fidelity of the original model at CFG 4.0 and 7.0, and is substantially better than the CFG 1.0 setting, which shows a marked degradation in fidelity.
Conclusion on trade-offs
Taken together, FairT2I appears to offer a favorable balance. It significantly enhances image diversity over standard CFG settings (4.0 and 7.0) and achieves diversity levels statistically comparable to the very low CFG 1.0 setting. While there is a statistically significant, albeit modest, reduction in CLIPScore compared to using CFG 4.0 and 7.0 with the original model, FairT2I maintains excellent image fidelity (best FID). This presents a notable advantage over simply lowering the CFG scale to 1.0 with the original model, which boosts diversity but at a considerable cost to image fidelity and without a clear advantage in text-image alignment over FairT2I. Therefore, FairT2I effectively increases diversity while largely preserving image quality and maintaining a competitive level of text-image alignment.
Table 15: Statistical comparison of FairT2I with Original Stable Diffusion (various CFG scales) using one-sided Mann-Whitney U tests. For all comparisons, the alternative hypothesis ($H_{A}$) tested is that the median score of Group 1 is stochastically greater than that of Group 2. Significance level $\alpha=0.05$.
Metric Group 1 Group 2 p-value Finding
BLIP2 Trace FairT2I CFG 7.0 <<0.000001 Group 1 significantly higher
FairT2I CFG 4.0 0.000002 Group 1 significantly higher
CFG 1.0 CFG 4.0 0.255153 Group 1 not significantly higher
CLIP Trace FairT2I CFG 7.0 <<0.000001 Group 1 significantly higher
FairT2I CFG 4.0 <<0.000001 Group 1 significantly higher
CFG 1.0 CFG 4.0 0.068690 Group 1 not significantly higher
CLIP Score CFG 7.0 FairT2I 0.020644 Group 1 significantly higher
CFG 4.0 FairT2I 0.008392 Group 1 significantly higher
CFG 1.0 FairT2I 0.664210 Group 1 not significantly higher
D.5 User Study Details
We conducted a user study using the crowdsourcing platform Prolific to collect annotations from a diverse group of participants. Prolific allows researchers to recruit participants according to predefined demographic quotas; in our case, we applied sex and ethnicity quotas to ensure balanced participation and obtain responses from a demographically varied pool of annotators.
All responses were collected via a Google Form. For each task, participants were presented with a text prompt and sets of images generated by four anonymized models (Model A, Model B, Model C, and Model D). They were asked to rate each model on a five-point Likert scale ($1=$ worst, $5=$ best) along three criteria: diversity, image quality, and text–image alignment. Each model output was displayed as a $3\times 3$ image grid aligned horizontally with white padding to facilitate visual comparison.
To promote consistent and reliable responses, the instruction section included concise explanations and rubrics defining each evaluation criterion. The complete task interface shown to participants is illustrated in Figure 11. Additionally, we enforced a regular-expression-based input format to ensure well-structured and standardized annotations across all responses.
Figure 11: User study interface provided to participants, including the task description, input prompts, generated images, and the multi-grid answer form for annotators.
We sampled 50 prompts from the Parti Prompts dataset [yu2022scaling], as detailed in subsection B.1. To keep the annotation workload manageable, these 50 prompts were divided into five separate tasks, each containing 10 prompts. Each task was assigned to a distinct group of participants to prevent repetition across annotators.
For each task, we recruited 20 independent annotators, resulting in a total of 100 participants overall. This configuration maintained an average completion time of approximately 20 minutes per participant. Each participant received an hourly rate of approximately £6.0, which is slightly below Prolific’s recommended fair payment rate but was deemed reasonable given the short task duration and minimal cognitive load.
D.6 Visual Results
This section presents additional visual comparisons for several input text prompts in Parti Prompt dataset. Figures 1215 illustrate images generated using our proposed FairT2I method alongside those generated with classifier-free guidance (CFG) at varying scales (7.0, 4.0, and 1.0). As noted in the captions, a guidance scale of 1.0 effectively corresponds to generation without the influence of classifier-free guidance. Our FairT2I method demonstrates an enhanced capability to diversify generation results, particularly for prompts that do not necessarily involve humans—spanning categories such as animals (Figures 12 and 15), objects (Figure 13), and abstract concepts (Figure 14). Each figure allows for a qualitative assessment of the different generation conditions for its respective prompt.
Figure 12: Generated images for the input text “Four deer surrounding a moose.” by classifier-free guidance (CFG) at guidance scales 7.0, 4.0, and 1.0 and FairT2I (Ours). A guidance scale of 1.0 corresponds to generation without CFG.
Figure 13: Generated images for the input text “a Ferrari Testarossa in front of the Kremlin” by classifier-free guidance (CFG) at guidance scales 7.0, 4.0, and 1.0 and FairT2I (Ours). A guidance scale of 1.0 corresponds to generation without CFG.
Figure 14: Generated images for the input text “meaning of life” by classifier-free guidance (CFG) at guidance scales 7.0, 4.0, and 1.0 and FairT2I (Ours). A guidance scale of 1.0 corresponds to generation without CFG.
Figure 15: Generated images for the input text “a cow” by classifier-free guidance (CFG) at guidance scales 7.0, 4.0, and 1.0 and FairT2I (Ours). A guidance scale of 1.0 corresponds to generation without CFG.